OpenStudy (anonymous):

How long does it take $10,000 to grow to $25,000 in an account with 3% simple interest

OpenStudy (anonymous):

i think the answer is 83.333

OpenStudy (anonymous):

r*p*n (.03)(10,000)*n=25000 solving for n we get 83.333

OpenStudy (anonymous):

you have to round up.

OpenStudy (anonymous):

the solution sheet my teacher gave me said the answer was 50...

OpenStudy (anonymous):

It should be the case that \(r>1\)

OpenStudy (luigi0210):

What formula are you using?

OpenStudy (anonymous):

r*p*n

OpenStudy (luigi0210):

Do you mean \[\LARGE P=A(1+r)^n\]

OpenStudy (anonymous):

Let \(r = 1.03\), not \(r=0.03\).

OpenStudy (anonymous):

I have no idea how they got 50 though.

OpenStudy (anonymous):

think it was a typo?

OpenStudy (anonymous):

what should the answer be?

OpenStudy (anonymous):

Okay I figured out how they did it.

OpenStudy (anonymous):

They used: \[ (1+r\times n)\times p \]

OpenStudy (anonymous):

Or, to make it much more simple for you, they used: \[ I = p\times r \times n \]But in this case, \(I = 25,000-10,000\) since \(I\) should not include initial principle.

OpenStudy (anonymous):

ive never seen that formula in my life.... -_-